OpenStudy (n8kd_da_beast):

What is the degree measure of 7π/10?

OpenStudy (n8kd_da_beast):

I do give medals to best answers.

OpenStudy (anonymous):

Just remember that π(radians)=180º so 7π/10 radians=(7/10)π=(7/10)*180º=126º conversely:126º=(126/180)*π=0.7π=(7/10)π radians

OpenStudy (n8kd_da_beast):

Thanks a lot. I really needed the help. :D

OpenStudy (alivejeremy):

180° is equivalent to π 126° is equivalent to 126/180 π Simplifying, we get 123/180 π = 7/10 π